OpenStudy (anonymous):

AWARDING MEDALS! Justify if completing the square is a good method for solving when the Discriminant is negative. Use any of your three functions as an example and respond in complete sentences.

OpenStudy (anonymous):

F(x) = (x-1) (x-2)

OpenStudy (anonymous):

please?

OpenStudy (campbell_st):

well completing the square will give you the answer in the same for as the general quadratic formula as an example consider \[x^2 +2x + 4 = 0\] the discriminant is \[\Delta = -12\] so using the general quadratic formula the solutions are \[x = \frac{-4 \pm \sqrt{2^2 - 4 \times 1 \times 4}}{2 \times 1}\] which simplifies to \[x = -1 \pm \sqrt{-3}...or ... x = -1 \pm i \sqrt{3}\] now if you rewrite the problem \[x^2 +2x = -4\] add halve the coefficient of the middle termed squared to both sides of the equation \[(\frac{2}{2})^2 = 1\] so you have \[x^2 + 2x + 1 = -3\] which is \[(x + 1)^2 = -3\] I'll let you finish to show that the solution is the same as using the general quadratic formula

OpenStudy (anonymous):

okay thank you!!